OpenStudy (pphalke):

nico is saving money for his college education. he invests some money at 6% and $1900 less than the amount at 5%. the investments produced a total of $191 interest in 1 yr. how much did he invest at each rate? please help i need steps in how to do it plzzzz i am begging

OpenStudy (tennis5518):

.05x+.04(x-1,900)=185 .05x+.04x-76=185 .09x=185+76 .09x=261 x=261/.09

OpenStudy (tennis5518):

x=2,900 amount invested @ 5%. 2,900-1,900=1,000 amount invested @ 4%. Proof: .05*2,900+.04*1,000=185 145+40=185 185=185

OpenStudy (tennis5518):

Let x be the amount he invests at 5%. Then the amount at 4% will be x - 1900. His interest for one year is 185 = 0.05x + 0.04(x - 1900) = 0.05x + 0.04x - 76

OpenStudy (tennis5518):

185 + 76 = 0.09x x = 261/0.09 = 2900 So, he invested $2900 at 5% and $1000 at 4%.

OpenStudy (pphalke):

its 6% and 5% not 4%